OpenStudy (anonymous):

How do I find 3/4 of 36?

OpenStudy (anonymous):

you should multiply them

OpenStudy (anonymous):

I know, But I need a method on how to explain it good to a 10 year old

OpenStudy (anonymous):

(3/4)(36)=(3/4)(4)(9)=(4/4)(3x9)=(1)(27)=27

OpenStudy (anonymous):

Find a quarter (divide 36 "counters" into 4 piles) and then count 3 piles....

OpenStudy (anonymous):

@hero How could you explain this to a 10-year-old boy?

OpenStudy (anonymous):

I know how to solve this myself, my brother is having trouble doing these kind of problems.

OpenStudy (anonymous):

And Im having trouble explaining how because to me it's obvious

hero (hero):

simply say some words translate to operations such as add, subtract, multiply and divide. For instance "and" means add. "is" means equal. And in this case, "of" means multiply. So 3/4 of 36 means 3/4 times 36.

OpenStudy (anonymous):

Or watch this video.I am sure you got a clue. http://www.khanacademy.org/math/arithmetic/fractions/v/multiplying-fractions

hero (hero):

Furthermore, 36 can also be written as a fraction because anything over 1 equals itself. So we can re-write it like so \[\frac{3}{4} \times \frac{36}{1}\] and if the 10 year old boy is already familiar with cancellation properties, you can tell him that it will become \[\frac{3}{1} \times \frac{9}{1} \] after cancellation

OpenStudy (anonymous):

Thank you all very much!